I need help trying to solve this problem.....I just don't understand how to set up the problem to solve. Thank you!
On a trip to take your brother to college, you and your parents travel 70 miles per hour on the way there and 55 miles per hour on the way home. The trip home took one hour more than the trip there. The total distance for the round trip was 805 miles. How many hours did it take for the whole trip?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

On a trip to take your brother to college, you and your parents travel 70 miles per hour on the way there and 55 miles per hour on the way home. The trip home took one hour more than the trip there. The total distance for the round trip was 805 miles. How many hours did it take for the whole trip?
---------------
805 miles RT --> 402.5 miles each way.
---
d = r*t
t = d/r
---
Going, t = 402.5/70 hours
Back, t = 402.5/55 hours
===============
That doesn't fit the data.
The difference is not 1 hour.

Answer by stanbon(75887)

On a trip to take your brother to college, you and your parents travel 70 miles per hour on the way there and 55 miles per hour on the way home. The trip home took one hour more than the trip there. The total distance for the round trip was 805 miles. How many hours did it take for the whole trip?
----
To-college DATA:
rate = 70 mph ; time = x hrs ; distance = r*t = 70x miles
-----
To-home DATA:
rate = 55 mph ; time = x+1 hrs ; distance = r*t = 55x + 55 miles
-------
Equation:
distance to + distance from = 805 miles
70x + 55x + 55 = 805 miles
125x = 750
x = 6 hrs (time to college)
x+1 = 7 hrs (time to home)
-----
Total time:: 13 hrs
==============
Cheers,
Stan H.
=========================